Cost per Unit
Variable costs
Direct material $940
Direct labor 600
Variable overhead 300
Fixed costs
Depreciation of equipment 500
Depreciation of building 275
Supervisory salaries 300
The company has an offer from Duvall Valves to produce the part for $2,000 per unit and supply 1,000 valves (the number needed in the coming year). If the company accepts this offer and shuts down production of valves, production workers and supervisors will be reassigned to other areas. The equipment cannot be used elsewhere in the company, and it has no market value. However, the space occupied by the production of the valve can be used by another production group that is currently leasing space for $55,000 per year.
What is the incremental savings of buying the valves? (The answer should be stated in a per-unit format and is a positive number)
Round to two decimal places.
Answer #1

Solution :-

Outsourcing Cost (2,000*1,000) = 2,000,000

Cost Savings

Direct Materials (940*1,000) = (940,000)

Direct Labor (600*1,000) = (600,000)

Variable Overhead (300*1,000) = (300,000)

Leased Space Savings = (55,000)

Supervisor Salary (300*1,000) = (300,000)

Incremental Savings = 195,000

Cost per unit = 195,000 / 1,000 = 195 per Unit